OpenStudy (anonymous):

What is 2/3x -5/6 = -1/3

OpenStudy (anonymous):

x = 3/4

OpenStudy (anonymous):

Make a common denominator for all of the fractions. In this case, 6. 4x/(6) - 5/6 = -2/6 Then, isolate the term containing a variable by adding 5/6 to both sides. 4x/6=3/6 Now multiply both sides by 6 and divide by 4

OpenStudy (anonymous):

and you'll get x = 3/4

OpenStudy (anonymous):

Thank You. Can someone show me the steps. You are great. 2/3x-5/6=-1/3

OpenStudy (anonymous):

Which step is confusing?